# library.bib

### v2

@@ 4.4 Graphical Interpretation
We construct a graph for each collection of set partitions $\sigma_{f},\tau_{f},\pi_{1},\dots,\pi_{l(w)}.$ This is similar to the construction of a surface from a matching datum by Magee in [Magee2021]. Indeed, the graph is essentially the same as the $1$–skeleton of the surfaces constructed there.
The graph we construct will be denoted $G\left(\sigma_{f},\tau_{f},\pi_{i}\right).$ The information in (24) will be encoded in the properties of $G\left(\sigma_{f},\tau_{f},\pi_{i}\right),$ allowing us to simplify further by analyzing the graph rather than dealing with tricky combinatorial arguments. Ultimately, from $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ we will construct a new graph from which we derive Theorem 1.5.
$G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ is a graph with $2kl(w)$ vertices, separated in to $k$ distinct subsets, each containing exactly $2l(w)$ vertices. We number the subsets from $1$ to $k$. For each $i\in\{1,\dots,k\}$, the vertices in subset $i$ are labeled
$\left(I_{f}^{1}\right)_{i},\dots\left(I_{f}^{|w|_{f}}\right)_{i},\left(J_{f}^{1}\right)_{i},\dots,\left(J_{f}^{|w|_{f}}\right)_{i}$ for each $f\in\{x_{1},\dots,x_{r}\}$.
Within each subset $i$, for each $f\in\{x_{1},\dots,x_{r}\}$ and for each $j\in\{1,\dots,|w|_{f}\},$ we draw a directed, $f$–labeled edge from the vertex labeled $\left(J_{f}^{j}\right)_{i}$ to the vertex labeled $\left(I_{f}^{j}\right)_{i}$. This gives a total of $kl(w)$ directed edges. We refer to the $i^{\mathrm{th}}$ subset of vertices as the $i^{\mathrm{th}}$ $w$–loop, see Figure 1.
@@
Figure 1: Here we have shown how to begin constructing $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ with $w=x_{1}x_{1}x_{2}x_{1}^{-1}x_{2}^{-1}$ and $k=2.$ There are $202kl(w)=$ 20 vertices, split in to $k=2$ distinct subsets, each containing $102l(w)=$ 10 vertices each. The vertices on the left (ie with the outer subscript “$1$”) are the $1^{\mathrm{{}^{\prime}st}}$ $w$–loop and the vertices on the right are the $2^{\mathrm{{}^{\prime}nd}}$ $w$–loop. There are $k|w|_{x_{1}}=6$ directed $x_{1}$–edges and $k|w|_{x_{2}}=4$ directed $x_{2}$–edges.
For each $f\in\{x_{1},\dots,x_{r}\}$, we add an undirected $\sigma_{f}$–edge between any two vertices $\left(J_{f}^{j}\right)_{i}$ and $\left(I_{f}^{j^{\prime}}\right)_{i^{\prime}}$ whenever $\sigma_{f}$ dictates that the corresponding indices must be equal in order for $\delta_{\sigma_{f}}\left(J_{f}^{1}\sqcup J_{f}^{2}\sqcup\dots\sqcup J_{f}^{|w|_{f}}\right)$ to be non–zero. This is illustrated in Figure 2.
@@
For each $f\in\{x_{1},\dots,x_{r}\}$, we add undirected $\tau_{f}$–edges similarly, illustrated in Figure 4.
Finally, for each $i\in\{1,\dots,l(w)\},$ we add an undirected $\pi_{i}$–edge between any two vertices whenever $\pi_{i}$ dictates that the corresponding indices must be equal in order for the corresponding inner product term in Definition 4.6 to be non–zero.
In short, for each collection of partitions $\sigma_{f},\tau_{f},\pi_{1},\dots,\pi_{l(w)},$ we draw an undirected edge between any two vertices for which the indices with the same label are necessarily equal for the conditions in Definition 4.6 to hold.
@@ 4.5 Obtaining the bound
In the case of $\sigma_{f}=\tau_{f}$ and with all $\pi_{1},\dots,\pi_{l(w)}$ corresponding to permutations (ie with $\sum_{i}\mathrm{del}(\pi_{i})=0$), we can apply [LouderWilton, Theorem 1.2] to obtain the bound of $n^{-k}.$ For the other cases, we need to prove a variant of this theorem that applies in our case. This section is dedicated to this task.
@@ 4.5.1 Stackings
Louder and Wilton [LouderWilton] first developed the notion of a stacking of a graph immersion and we adapt their definitions slightly to suit our case.
@@
Let $\Gamma$ be a finite graph and $\mathbb{S}$ a $1$–complex with an immersion $\Lambda:\mathbb{S}\to\Gamma.$ A stacking is an embedding $\hat{\Lambda}:\mathbb{S}\to\Gamma\times\mathbb{R}$ such that $\pi\hat{\Lambda}=\Lambda,$ where $\pi:\Gamma\times\mathbb{R}\to\Gamma$ is the trivial $\mathbb{R}$–bundle.
@@ 4.5.2 Bounding expected character
By Remark 1.6, we only need to prove the bound in Theorem 1.5 in the case where $w$ is not the identity or a proper power. From now, assume that $w$ is not a proper power.
For each $\sigma_{f},\tau_{f},\pi_{i},$ construct the graph $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ from Section 4.4, consisting of $kl(w)$ disjoint, directed $f$ edges organised into $k$ $w$–loops and then marking on the blocks of the partitions $\sigma_{f},\tau_{f},\pi_{i}$ using undirected edges.
@@
  • gluing together any vertices that are connected by a $\sigma_{f},\tau_{f}$ or $\pi_{i}$–edge,
  • if $\sigma_{f}$ and $\tau_{f}$ connect both the initial and terminal vertices of some collection of $f$–edges, we merge these in to a single $f$–edge, see figure 5.
@@
Every block of $\sigma_{f}\wedge\tau_{f}$ of size $p$ corresponds to some collection of $f$–edges in $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ of size $p$, whose initial and terminal vertices have been glued together and whose edges have been merged in the construction of $\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right).$ Every $f$–edge in $\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ then corresponds to a block of $\sigma_{f}\wedge\tau_{f}.$
@@
and seek to apply [LouderWilton, Theorem 1.2] directly. Unfortunately, in gluing the vertices of $\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ together, one may lose the distinctness property that no two vertices of $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ labeled, say, $\left(J_{f}^{i}\right)_{j_{1}}$ and $\left(J_{f}^{i}\right)_{j_{2}}$ with $j_{1}\neq j_{2}$, can be glued together, which would then weaken the bound achieved by applying [LouderWilton, Theorem 1.2]. Therefore, we proceed with our extension of their result.
Define $\mathcal{B}_{r}\overset{\mathrm{def}}{=}\bigvee_{i=1}^{r}S_{i}^{1},$ the bouquet of $r$ oriented circles, labeled $x_{1},\dots,x_{r}$ and with wedge point labeled $o$. Define $W$ to be the graph consisting of a single cycle of oriented $f$–labeled edges that, when traversed in the correct direction, reads out the word $w.$ Then $W$ comes equipped with an obvious primitive (since $w$ is not a proper power) immersion
@@
By [LouderWilton, Lemma 3.4], this immersion has a stacking
@@
We denote by $\mathbb{S}$ the components of the fibre product graph defined by following the paths in $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ which alternate between $f$–edges and $\pi_{i}$–edges. This defines some partition of $w^{k}.$
The connected components of $\mathbb{S}$ are either circles that read out $w^{k_{i}}$ or pieces, which consist of two closed endpoints and a chain of vertices of valence two in between them. Each piece reads out some subword of $w^{k}.$ The concatenation of the words spelled out by all the connected components of $\mathbb{S}$ is exactly $w^{k}$ and the number of pieces of $\mathbb{S}$ is exactly $\sum_{i}\mathrm{del}\left(\pi_{i}\right).$
@@
By the same argument as [LouderWilton, Lemma 2.5], we have a stacking,
@@
  • a closed arc: a connected and simply connected union of vertices and interiors of edges, with both ends closed (ie, a piece).
@@
then the following lemma is immediate. [^4] We are especially indebted to Noam Ta Shma for pointing out an error in our original argument for proving the extension of Wise’s $w$–cycles conjecture and for his feedback on the correction presented from this point onward.
@@
Neither $\mathcal{A}_{\hat{\Lambda}^{\prime}}$ or $\mathcal{B}_{\hat{\Lambda}^{\prime}}$ contain a circle, a closed arc reading out a word of length $\geq l(w)$ or an open/half–open arc reading out a word of length $>l(w).$
@@
Let $C\subseteq\mathbb{S}$ be a circle, a closed arc reading out a word of length $\geq l(w)$, or an open/half–open arc reading out a word of length $>l(w),$ which is contained inside $\mathcal{A}_{\hat{\Lambda}^{\prime}}$. Then the map
@@
  • closed arcs that read out words of length $<l(w).$
@@
$\mathcal{A}\overset{\mathrm{def}}{=}\#\big{\{}\mathrm{ends\ of\ connected\ components\ of}\ \mathcal{A}_{\hat{\Lambda}^{\prime}}\big{\}}$
@@
=$\mathcal{B}\overset{\mathrm{def}}{=}\#\big{\{}\mathrm{ends\ of\ connected\ components\ of}\ \mathcal{B}_{\hat{\Lambda}^{\prime}}\big{\}}.$
@@
We have $\mathcal{A}\geq 2k$ and $\mathcal{B}\geq 2k.$
@@
Pick any vertex $w\in W$ such that $w\in\mathcal{A}_{\hat{\Lambda}}.$ Then $w$ has at least $k$ preimages $\delta^{-1}(w)$ in $\mathbb{S},$ all of which must be contained in $\mathcal{A}_{\hat{\Lambda}^{\prime}}.$ If two such preimages belong to the same connected component $C$ of $\mathcal{A}_{\hat{\Lambda}^{\prime}},$ then $C$ must either be a circle, a closed arc reading out a word of length $\geq l(w)$ or an open/half–open arc reading out a word of length $>l(w),$ contradicting Lemma 4.17. Thus, there must be at least $k$ connected components in $\mathcal{A}_{\hat{\Lambda}^{\prime}},$ so there must be at least $2k$ ends.
Repeating the argument with $w\in\mathcal{B}_{\hat{\Lambda}}$ proves the lemma for $\mathcal{B}.$
@@
$\mathcal{A}=-2\chi+2\mathcal{A}_{\mathrm{closed}}$
@@
$\mathcal{\mathcal{B}}=-2\chi+2\mathcal{B}_{\mathrm{closed}}.$
@@
since the total number of closed ends in $\mathbb{S}$ is $2\sum_{i}\mathrm{del}\left(\pi_{i}\right)$ and each closed end in $\mathbb{S}$ can be, at most, a closed end in $\mathcal{A}_{\hat{\Lambda}^{\prime}}$ or a closed end in $\mathcal{B}_{\hat{\Lambda}^{\prime}},$ but not both, since that would imply that its image in $\Gamma$ is covered only once by the immersion $\Lambda^{\prime}:\mathbb{S}\to\Gamma$, contradicting Lemma 4.15.
Combining (26) with (25) proves Theorem 1.5.
@@
$=$ $d_{\lambda}\sum_{\sigma_{f},\tau_{f}}^{\star}\sum_{\pi_{1},\dots,\pi_{l(w)}}^{\leq S_{k}}\left(\prod_{i=1}^{l(w)}c(n,k,\lambda,\pi_{i})\right)\left(\prod_{f\in\{x_{1},\dots,x_{m}\}}\mathrm{Wg}_{n,\left(|w|_{f}k\right)}\left(\sigma_{f},\tau_{f}\right)\right)$
@@
For each $f\in\{x_{1},\dots,x_{m}\},$ for each $\sigma_{f}\in\overset{\star}{\mathrm{Part}}\left(\left[|w|_{f}k\right]\right),$
@@
$c(n,k,\lambda,\pi)=\frac{(-1)^{k}d_{\lambda}^{2}(n)_{2k}}{\left(k!\right)^{2}(n)_{\lambda}}\left[\sum_{\sigma,\sigma^{\prime}\in S_{k}}\sum_{\pi^{\prime}\leq\iota\left(\sigma^{-1}\sigma^{\prime}\right)\wedge\pi}\frac{\mu\left(\pi^{\prime},\pi\right)(-1)^{|\pi^{\prime}|}\chi^{\lambda}\left(\sigma\right)\chi^{\lambda}\left(\sigma^{\prime}\right)}{(n)_{|\pi^{\prime}|}}\right],$
where $(n)_{\lambda}=\prod_{j=1}^{k}\left(n-k+1+\check{\lambda}_{j}-j\right).$ The term inside the square bracket is equal to
$\frac{\underset{\sigma,\sigma^{\prime}\in S_{k}}{\sum}\chi^{\lambda}(\sigma)\chi^{\lambda}\left(\sigma^{\prime}\right)+\dots+\kappa_{\pi}\left(n-|\pi|\right)\dots\left(n-2k+1\right)}{(n)_{2k}},$
with
$\kappa_{\pi}=\left(\underset{\sigma,\sigma^{\prime}\ \mathrm{s.t.}\ \pi\leq\iota\left(\sigma^{-1}\sigma^{\prime}\right)}{\sum}(-1)^{|\pi|}\chi^{\lambda}\left(\sigma\right)\chi^{\lambda}\left(\sigma^{\prime}\right)\right).$
So then $c(n,k,\lambda,\pi)$ is equal to
$\frac{h_{k,\lambda,\pi}(n)}{(n)_{\lambda}},$
@@
$h_{k,\lambda,\pi}=\frac{(-1)^{k}d_{\lambda}^{2}}{\left(k!\right)^{2}}\Bigg{[}\underset{\sigma,\sigma^{\prime}\in S_{k}}{\sum}\chi^{\lambda}(\sigma)\chi^{\lambda}\left(\sigma^{\prime}\right)+\kappa_{\pi}\left(n-|\pi|\right)\dots\left(n-2k+1\right)\Bigg{]},$
a polynomial in $n$ of maximum degree $2k-|\pi|=k-\mathrm{del}(\pi).$
It follows that
$\prod_{i=1}^{l(w)}c(n,k,\lambda,\pi_{i})=\frac{\prod_{i=1}^{l(w)}h_{k,\lambda,\pi_{i}}(n)}{\left(n\right)_{\lambda}^{l(w)}}.$
Hence,
@@
$p_{k,\lambda,w}\left(n\right)=d_{\lambda}\sum_{\sigma_{f}}^{\star}\sum_{\pi_{1},\dots,\pi_{l(w)}}^{\leq S_{k}}\left(\prod_{f}g_{\sigma_{f},\tau_{f}}(n)\right)\left(\prod_{i=1}^{l(w)}h_{k,\lambda,\pi_{i}}(n)\right)\mathcal{N}\left(\sigma_{f},\tau_{f},\pi_{i}\right).$
@@
    $\sum_{i=1}^{l(w)}k-\mathrm{del}(\pi_{i})=kl(w)-\sum_{i}\mathrm{del(\pi_{i})},$
  • $\mathcal{N}\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ has maximum degree
    $-k+\sum_{i}\mathrm{del}(\pi_{i})+\sum_{f}|\sigma_{f}\wedge\tau_{f}|,$
@@
$=$ $n^{kl(w)}\prod_{d=1}^{m}\prod_{c=1}^{|w|_{x_{d}}k-1}\left(1-c.\frac{1}{n}\right).$
@@
$\left(n\right)_{\lambda}^{l(w)}=n^{kl(w)}\left[\prod_{j=1}^{k}\left(1+\left(1+\check{\lambda}_{j}-j-k\right).\frac{1}{n}\right)\right]^{l(w)}.$
@@
$\hat{g}(x)=\prod_{c=1}^{q}\left(1-cx\right)^{q}\left[\prod_{j=1}^{k}\left(1+\left(1+\check{\lambda}_{j}-j-k\right)x\right)\right]^{q}.$
@@
$\hat{g}(x)\Biggr{|}\prod_{c=1}^{q}\left(1-cx\right)^{q}\left[\prod_{j=1}^{2k}\left(1+\left(1-j\right)x\right)\right]^{q}=g_{q,k}(x).$
@@
[…]

### v3

@@ 4.4 Graphical Interpretation
We construct a graph for each collection of set partitions $\sigma_{f},\tau_{f},\pi_{1},\dots,\pi_{l(w)}.$ This is similar to the construction of a surface from a matching datum by Magee in [Mag25]. Indeed, the graph is essentially the same as the 11–skeleton of the surfaces constructed there.
The graph we construct will be denoted $G\left(\sigma_{f},\tau_{f},\pi_{i}\right).$ The information in (23) will be encoded in the properties of $G\left(\sigma_{f},\tau_{f},\pi_{i}\right),$ allowing us to simplify further by analyzing the graph rather than dealing with tricky combinatorial arguments. Ultimately, from $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ we will construct a new graph, from which we derive Theorem 1.6.
$G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ is a graph with $2kl(w)$ vertices, separated in to $k$ distinct subsets, each containing exactly $2l(w)$ vertices. We number the subsets from $1$ to $k$. For each $i\in\{1,\dots,k\}$, the vertices in subset $i$ are labeled $\left(I_{f}^{1}\right)_{i},\dots\left(I_{f}^{|w|_{f}}\right)_{i},\left(J_{f}^{1}\right)_{i},\dots,\left(J_{f}^{|w|_{f}}\right)_{i}$ for each $f\in\{x_{1},\dots,x_{r}\}$.
Within each subset $i$, for each $f\in\{x_{1},\dots,x_{r}\}$ and for each $j\in\{1,\dots,|w|_{f}\},$ we draw a directed, ff–labeled edge from the vertex labeled $\left(J_{f}^{j}\right)_{i}$ to the vertex labeled $\left(I_{f}^{j}\right)_{i}$. This gives a total of $kl(w)$ directed edges. We refer to the $i^{\mathrm{th}}$ subset of vertices as the $i^{\mathrm{th}}$ ww–loop, see Figure 1.
@@
Figure 1: Here we have shown how to begin constructing $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ with $w=x_{1}x_{1}x_{2}x_{1}^{-1}x_{2}^{-1}$ and $k=2.$ There are $2kl(w)=20$ vertices, split in to $k=2$ distinct subsets, each containing $2l(w)=10$ vertices each. The vertices on the left (i.e. with the outer subscript “11”) are the $1^{\mathrm{st}}$ ww–loop and the vertices on the right are the $2^{\mathrm{nd}}$ ww–loop. There are $k|w|_{x_{1}}=6$ directed $x_{1}$–edges and $k|w|_{x_{2}}=4$ directed $x_{2}$–edges.
For each $f\in\{x_{1},\dots,x_{r}\}$, we add an undirected $\sigma_{f}$–edge between any two vertices $\left(J_{f}^{j}\right)_{i}$ and $\left(J_{f}^{j^{\prime}}\right)_{i^{\prime}}$ whenever $\sigma_{f}$ dictates that the corresponding indices must be equal in order for $\delta_{\sigma_{f}}\left(J_{f}^{1}\sqcup J_{f}^{2}\sqcup\dots\sqcup J_{f}^{|w|_{f}}\right)$ to be non–zero. This is illustrated in Figure 2.
@@
For each $f\in\{x_{1},\dots,x_{r}\}$, we add undirected $\tau_{f}$–edges similarly, illustrated in Figure 4. Finally, for each $i\in\{1,\dots,l(w)\},$ we add an undirected $\pi_{i}$–edge between any two vertices whenever $\pi_{i}$ dictates that the corresponding indices must be equal in order for the corresponding inner product term in Definition 4.5 to be non–zero.
In short, for each collection of set partitions $\sigma_{f},\tau_{f},\pi_{1},\dots,\pi_{l(w)},$ we draw an undirected edge between any two vertices for which the indices with the same label are necessarily equal for the conditions in Definition 4.5 to hold.
@@ 4.5 Obtaining the bound
With all $\pi_{1},\dots,\pi_{l(w)}$ corresponding to permutations (i.e. with $\sum_{i}\mathrm{del}(\pi_{i})=0$), we can apply [LW17, Theorem 1.2] after (24) below to obtain the bound of $n^{-k}.$ For the other cases, we need to prove a variant of this theorem that applies in our case. This section is dedicated to this task.
@@ 4.5.1 Stackings
Louder and Wilton [LW17] first developed the notion of a stacking of a graph immersion and we adapt their definitions slightly to suit our case.
@@
Let $\Gamma$ be a finite graph and $\mathbb{S}$ a 11–complex with an immersion $\Lambda:\mathbb{S}\to\Gamma.$ A stacking is an embedding $\hat{\Lambda}:\mathbb{S}\to\Gamma\times\mathbb{R}$ such that $\pi\hat{\Lambda}=\Lambda,$ where $\pi:\Gamma\times\mathbb{R}\to\Gamma$ is the trivial $\mathbb{R}$–bundle.
@@ 4.5.2 Bounding expected character
By Remark 1.7, we only need to prove the bound in Theorem 1.6 in the case where $w$ is not a proper power. From now, assume that $w$ is not a proper power.
For each $\sigma_{f},\tau_{f},\pi_{i},$ construct the graph $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ from Section 4.4, consisting of $kl(w)$ disjoint, directed $f$ edges organised into $k$ ww–loops and then marking on the blocks of the partitions $\sigma_{f},\tau_{f},\pi_{i}$ using undirected edges.
@@
  • gluing together any vertices that are connected by a $\sigma_{f},\tau_{f}$ or $\pi_{i}$–edge (and then deleting the $\sigma_{f},\tau_{f}$ or $\pi_{i}$–edges),
  • if $\sigma_{f}$ and $\tau_{f}$ connect both the initial and terminal vertices of some collection of ff–edges, we merge these into a single ff–edge, see Figure 5.
Definition 4.10.
@@
Every block of $\sigma_{f}\wedge\tau_{f}$ of size $p$ corresponds to some collection of ff–edges in $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ of size $p$, whose initial and terminal vertices have been glued together and whose edges have been merged in the construction of $\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right).$ Every ff–edge in $\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ then corresponds to a block of $\sigma_{f}\wedge\tau_{f}.$
@@
and seek to apply [LW17, Theorem 1.2] directly. Unfortunately, in gluing the vertices of $\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ together, one may lose the distinctness property that no two vertices of $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ labeled, say, $\left(J_{f}^{i}\right)_{j_{1}}$ and $\left(J_{f}^{i}\right)_{j_{2}}$ with $j_{1}\neq j_{2}$, can be glued together, which would then weaken the bound achieved by applying [LW17, Theorem 1.2]. Therefore, we proceed with our extension of their result.
Define $\mathcal{B}_{r}\overset{\mathrm{def}}{=}\bigvee_{i=1}^{r}S_{i}^{1},$ the bouquet of $r$ oriented circles, labeled $x_{1},\dots,x_{r}$ and with wedge point labeled $o$. Define $W$ to be the graph consisting of a single cycle of oriented ff–labeled edges that, when traversed in an appropriate direction from an appropriate vertex, reads out the word $w.$ Then $W$ comes equipped with an obvious primitive (since $w$ is not a proper power) immersion
@@
By [LW17, Lemma 3.4], this immersion has a stacking
@@
We denote by $\mathbb{S}$ the components of the fibre product graph defined by following the paths in $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ which alternate between ff–edges and $\pi_{i}$–edges. This defines some partition of $w^{k}.$ The connected components of $\mathbb{S}$ are either circles that read out $w^{k_{i}}$ or pieces, which consist of two closed endpoints and a chain of vertices of valence two in between them. Each piece reads out some subword of $w^{k}.$ The concatenation of the words spelled out by all the connected components of $\mathbb{S}$ is exactly $w^{k}$ and the number of pieces of $\mathbb{S}$ is exactly $\sum_{i}\mathrm{del}\left(\pi_{i}\right).$
@@
By the same argument as [LW17, Lemma 2.5], we have a stacking,
@@
  • a closed arc: a connected and simply connected union of vertices and interiors of edges, with both ends closed (i.e. a piece).
@@
then the following lemma is immediate. [^5] We are especially indebted to Noam Ta Shma for pointing out an error in our original argument for proving the extension of Wise’s ww–cycles conjecture and for his feedback on the correction presented from this point onward.
@@
Neither $\mathcal{A}_{\hat{\Lambda}^{\prime}}$ or $\mathcal{B}_{\hat{\Lambda}^{\prime}}$ contain a circle, a closed arc reading out a word of length $\geq l(w)$ or an open/half–open arc reading out a word of length >l(w).>l(w).
@@
Let $C\subseteq\mathbb{S}$ be a circle, a closed arc reading out a word of length $\geq l(w)$, or an open/half–open arc reading out a word of length >l(w),>l(w), which is contained inside $\mathcal{A}_{\hat{\Lambda}^{\prime}}$. Then the map
@@
  • closed arcs that read out words of length <l(w).<l(w).
@@
$\mathcal{A}_{\mathrm{end}}\overset{\mathrm{def}}{=}\#\big\{\mathrm{ends\ of\ connected\ components\ of}\ \mathcal{A}_{\hat{\Lambda}^{\prime}}\big\}$
@@
$\mathcal{B}_{\mathrm{end}}\overset{\mathrm{def}}{=}\#\big\{\mathrm{ends\ of\ connected\ components\ of}\ \mathcal{B}_{\hat{\Lambda}^{\prime}}\big\}.$
@@
We have $\mathcal{\mathcal{A}_{\mathrm{end}}}\geq 2k$ and $\mathcal{B}_{\mathrm{end}}\geq 2k.$
@@
Pick any vertex $w\in W$ such that $w\in\mathcal{A}_{\hat{\Lambda}}$. Then $w$ has at least $k$ preimages $\delta^{-1}(w)$ in $\mathbb{S},$ all of which must be contained in $\mathcal{A}_{\hat{\Lambda}^{\prime}}.$ If two such preimages belong to the same connected component $C$ of $\mathcal{A}_{\hat{\Lambda}^{\prime}},$ then $C$ must either be a circle, a closed arc reading out a word of length $\geq l(w)$ or an open/half–open arc reading out a word of length >l(w),>l(w), contradicting Lemma 4.16. Thus, there must be at least $k$ connected components in $\mathcal{A}_{\hat{\Lambda}^{\prime}},$ so there must be at least $2k$ ends. Repeating the argument with $w\in\mathcal{B}_{\hat{\Lambda}}$ proves the lemma for $\mathcal{B}_{\mathrm{end}}.$
@@
$\mathcal{A}_{\mathrm{end}}=-2\chi+2\mathcal{A}_{\mathrm{closed}}$
@@
$\mathcal{\mathcal{B}_{\mathrm{end}}}=-2\chi+2\mathcal{B}_{\mathrm{closed}}.$
@@
since the total number of closed ends in $\mathbb{S}$ is $2\sum_{i}\mathrm{del}\left(\pi_{i}\right)$ and each closed end in $\mathbb{S}$ can be, at most, a closed end in $\mathcal{A}_{\hat{\Lambda}^{\prime}}$ or a closed end in $\mathcal{B}_{\hat{\Lambda}^{\prime}},$ but not both, since that would imply that its image in $\Gamma$ is covered only once by the immersion $\Lambda^{\prime}:\mathbb{S}\to\Gamma$, contradicting Lemma 4.14. Combining (25) with (24) proves Theorem 1.6.
@@
$=$ $\frac{1}{d_{\lambda}}\sum_{\sigma_{f},\tau_{f}}^{\star}\sum_{\pi_{1},\dots,\pi_{l(w)}}^{\leq S_{k}}\left(\prod_{i=1}^{l(w)}c(n,k,\lambda,\pi_{i})\right)\left(\prod_{f\in\{x_{1},\dots,x_{m}\}}\mathrm{Wg}_{n,\left(|w|_{f}k\right)}\left(\sigma_{f},\tau_{f}\right)\right)$
@@
For each $f\in\{x_{1},\dots,x_{m}\},$ for each $\sigma_{f},\tau_{f}\in\overset{\star}{\mathrm{Part}}\left(\left[|w|_{f}k\right]\right),$
@@
$c(n,k,\lambda,\pi)=\frac{d_{\lambda^{+}(n)}(-1)^{|\pi|+k}}{(n)_{|\pi|}}\sum_{\overset{\tau\in S_{k}}{{\scriptstyle{\scriptscriptstyle\iota(\tau)\geq\pi}}}}\chi^{\lambda}(\tau).$
Using the hook–length formula, this is equal to
$\frac{\left[\frac{(-1)^{k+|\pi|}}{k!}\sum_{\overset{\tau\in S_{k}}{{\scriptstyle{\scriptscriptstyle\iota(\tau)\geq\pi}}}}\chi^{\lambda}(\tau)\right](n-|\pi|)(n-|\pi|-1)\dots(n-2k+1)}{(n)_{\lambda}},$
where $(n)_{\lambda}=\prod_{j=1}^{k}\left(n-k+1-\check{\lambda}_{j}-j\right).$ It follows that
$\prod_{i=1}^{l(w)}c(n,k,\lambda,\pi_{i})=\frac{\prod_{i=1}^{l(w)}h_{k,\lambda,\pi_{i}}(n)}{(n)_{\lambda}^{l(w)}},$
@@
$h_{k,\lambda,\pi_{i}}(n)=\left[\frac{(-1)^{k+|\pi|}}{k!}\sum_{\overset{\tau\in S_{k}}{{\scriptstyle{\scriptscriptstyle\iota(\tau)\geq\pi}}}}\chi^{\lambda}(\tau)\right](n-|\pi|)(n-|\pi|-1)\dots(n-2k+1)$
is a polynomial in $n$ of degree $2k-|\pi_{i}|=k-\mathrm{del}\left(\pi_{i}\right).$ Hence,
@@
$p_{k,\lambda,w}\left(n\right)=\frac{1}{d_{\lambda}}\sum_{\sigma_{f}}^{\star}\sum_{\pi_{1},\dots,\pi_{l(w)}}^{\leq S_{k}}\left(\prod_{f}g_{\sigma_{f},\tau_{f}}(n)\right)\left(\prod_{i=1}^{l(w)}h_{k,\lambda,\pi_{i}}(n)\right)\mathcal{N}\left(\sigma_{f},\tau_{f},\pi_{i}\right).$
@@
    $\sum_{i=1}^{l(w)}\left(k-\mathrm{del}(\pi_{i})\right)=kl(w)-\sum_{i}\mathrm{del(\pi_{i})},$
  • $\mathcal{N}\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ has maximum degree [^6] This follows from the simple observation that $\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ has Euler characteristic $\leq\sum_{i}\mathrm{del}(\pi_{i})$, regardless of if $w$ is a proper power or not.
    $\sum_{i}\mathrm{del}(\pi_{i})+\sum_{f}|\sigma_{f}\wedge\tau_{f}|,$
@@
$=$ $n^{kl(w)}\prod_{d=1}^{m}\prod_{c=1}^{|w|_{x_{d}}k-1}\left(1-c\frac{1}{n}\right).$
@@
$\left(n\right)_{\lambda}^{l(w)}=n^{kl(w)}\left[\prod_{j=1}^{k}\left(1+\left(1+\check{\lambda}_{j}-j-k\right)\frac{1}{n}\right)\right]^{l(w)}.$
@@
$\hat{g}(x)=\prod_{c=1}^{kq}\left(1-cx\right)^{q}\left[\prod_{j=1}^{k}\left(1+\left(1+\check{\lambda}_{j}-j-k\right)x\right)\right]^{q}.$
@@
$\hat{g}(x)\Biggr|\prod_{c=1}^{kq}\left(1-cx\right)^{q}\left[\prod_{j=1}^{2k}\left(1+\left(1-j\right)x\right)\right]^{q}=g_{q,k}(x).$
@@
$\begin{aligned} &P_{w,k,\lambda}\left(\frac{1}{n}\right)\\ =&\hat{P}_{w,k,\lambda}\left(\frac{1}{n}\right)\frac{g_{q,k}\left(\frac{1}{n}\right)}{\tilde{g}\left(\frac{1}{n}\right)}\end{aligned}.$